Graph the equation x+3=y by finding three points on the line
Answer by Cromlix(4381)

Substitute x values into your equation to obtain the y coordinate.
x= -2 => -2 + 3 = 1 : y = 1
x = 0 => 0 + 3 = 3 : y = 3
x = 2 => 2 + 3 = 5 : y = 5
So your 3 points on this line are:
(-2, 1) (0, 3) and (2, 5)